# An open catalogue of digital world literature with likes

### v1

@@ 3.3 Release
Additionally, $74.5\%$ of the catalogue features a list of Genres and $98.8\%$ includes the number of reviews.
@@
The feature coverage for these datasets is identical to the primary English corpus, with the sole exception of Genres, which falls to $60.9\%$ for German, $55.2\%$ for Spanish, and $50.4\%$ for French. We speculate this is not an artefact of our matching but rather reflects a more sparse metadata structure for non-Anglophone works within Goodreads itself.
@@
  • The formatted metadata extracted for every Work, Edition, and Author from Goodreads.
@@ 4 Legal and ethical considerations
The datasets generated during and/or analysed during the current study will be made available in a public data repository.
To facilitate the diachronic analysis of these datasets, our primary English and secondary language corpora were added to Gallicagram [azoulay2021gallicagram], an n-gram viewer designed to explore term frequencies across its collection of French and international datasets [^6] https://gallicagram-react.vercel.app/. Gallicagram also enables access to its n-gram data via its api.

### v2

@@ 3.3 Release
Additionally, $84.04\%$ of the catalogue features a list of Genres and $99.23\%$ includes the number of reviews.
@@
The feature coverage for these datasets is identical to the primary English corpus, with the sole exception of Genres, which falls to $77.4\%$ for Spanish, $75.5\%$ for German, and $70.1\%$ for French. We speculate this is not an artefact of our matching but rather reflects a more sparse metadata structure for non-Anglophone works within Goodreads itself.
@@
  • The formatted metadata extracted for every Work, Edition, and Author from the Goodreads crawl.
@@ 4 Legal and ethical considerations
The datasets generated and/or analysed during the present study are available in the Zenodo public data repository: doi.org/10.5281/zenodo.17609566.
The detailed description of the datasets’ structure and metadata schemas is available on GitHub: github.com/mazieres/MajinBook.
To facilitate the diachronic analysis of these datasets, our primary English and secondary language corpora were added to Gallicagram [^6] https://gallicagram-react.vercel.app/, an n-gram viewer designed to explore term frequencies across its collection of French and international datasets [azoulay2021gallicagram]. Gallicagram also enables access to its n-gram data via its api.